### v2

# Quantum Non-Local Magic

Abstract
Quantum entanglement and quantum magic are fundamental resources that characterize distinct aspects of a quantum state: entanglement reflects non-local correlations, while magic quantifies the deviation from stabilizer states. A quantum state becomes a valuable resource for applications like universal quantum computation only when both quantities are present. Here, we propose that quantum non-local magic (NM) serves as an effective measure of this combined resource, incorporating both entanglement and magic. We demonstrate that NM can be precisely computed for two-qubit pure states, where it is directly related to the entanglement spectrum. We then extend the definition of NM to mixed states and explore its presence in many-body quantum systems, revealing that the two-point NM decays according to a power law in critical states. Furthermore, we explore measurement-induced NM and uncover an intriguing phenomenon termed “magic swapping”, analogous to entanglement swapping, wherein post-measurement NM decays more slowly than any pre-measurement correlations. Our results thus represent a pivotal step towards accurately quantifying the “quantumness” of a state and reveal the potential for manipulating this resource through measurements.
Quantum entanglement, which captures the inability of a quantum state to be decomposed into product states of its subsystems, is a fundamental manifestation of quantum mechanics [1]. It serves as a cornerstone for numerous applications, including quantum computation, quantum simulation and quantum communication [2, 3]. Entanglement is conventionally viewed as a non-local resource that cannot be generated through local operations alone, and is typically quantified by the von Neumann entropy for pure states or the entanglement negativity for mixed states [4, 5, 6, 7, 8]. Since an entangled state of $N$ qubits generally requires storing $O(2^{N})$ complex amplitudes for its classical representation, quantum algorithms leveraging such states can achieve exponential speedups over direct classical simulations, potentially even surpassing any known classical algorithm—a concept referred to as quantum supremacy [9, 10, 11].
Nevertheless, entanglement alone does not fully capture the “quantumness” of a state. Certain highly entangled states, known as stabilizer states, can still be efficiently described and simulated [12, 13, 14]. A stabilizer state is uniquely specified by a set of stabilizer generators, which requires only $O(N^{2})$ memory space. Furthermore, when a stabilizer state evolves under unitary gates from the Clifford group or under Pauli measurements, it remains within the stabilizer formalism, allowing for efficient classical simulation of the entire process. Consequently, non-stabilizerness, often referred to as “quantum magic”, is equally crucial for a quantum state to serve as a valuable resource in demonstrating quantum supremacy [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. It is worth emphasizing that a key difference between these two resources is that entanglement remains invariant under local unitary transformations, whereas quantum magic is basis-dependent, implying that it can be altered through local operations.
A natural question thus arises: Is there a suitable measure of the genuinely non-classical resources that integrates both entanglement and magic? Moreover, is this measure computationally tractable? Theoretically, such a measure would deepen our understanding of the connection between these two seemingly distinct properties [32, 33], while practically, it would help characterize algorithmic complexity and guide the selection of initial states in quantum computations and simulations. With the existence of such a measure, it becomes pertinent to identify resourceful states [34]. Furthermore, a more intriguing question is whether this resource can be manipulated through measurement.
In this Letter, we address these questions by investigating quantum non-local magic (NM), which quantifies the minimal amount of magic contained in a bipartite state after optimizing over all possible local unitary transformations on each subsystem [35, 36, 37]. We first demonstrate that NM is a promising candidate for characterizing quantum resources and is applicable to any faithful and additive magic measure. Although previous studies have focused on providing bounds and estimates of NM, we show that NM is analytically tractable for a two-qubit pure state and directly related to the entanglement spectrum. We then investigate which states exhibit non-zero NM. Our results indicate that NM is ubiquitous in typical two-qubit pure states, with Haar-random states having the highest probability of attaining maximal NM. Extending the definition to mixed state, we further examine the behavior of two-point NM in critical many-body states, using both the transverse-field Ising model (TFIM) at the critical point and a monitored Haar-random unitary circuit (MHC) with critical measurement rate as representative examples [38, 39, 40]. We find that two-point NM decays following a power law, mirroring the behavior of mutual information. Finally, we explore whether NM can be manipulated through measurements, akin to the manipulation of entanglement. Intriguingly, measurement-induced NM in a critical MHC decays more slowly than any correlation present prior to measurement, revealing a “magic swapping” phenomenon reminiscent of entanglement swapping [41, 42].
Non-local magic. We consider a bipartite pure state $\left|\Psi_{AB}\right\rangle$, where $A$ and $B$ represent two subsystems. As illustrated in Fig. 1(a), the NM is defined as [35, 36]:
$\mathcal{M}_{AB}(\left|\Psi_{AB}\right\rangle)=\min_{U_{A}\otimes U_{B}}M(U_{A}\otimes U_{B}\left|\Psi_{AB}\right\rangle),$ (1)
where $M$ is a quantum magic measure, and $U_{A}$ and $U_{B}$ are unitary gates acting on subsystems $A$ and $B$, respectively. We will omit the subscript of $\mathcal{M}_{AB}$ when the bipartition is clear from the context. In operational terms, a finite NM indicates that the state cannot be efficiently represented as a stabilizer state, regardless of the local unitaries applied to subsystems $A$ and $B$ in an effort to reduce the magic. By definition, $\mathcal{M}$ is symmetric under the exchange of subsystems $A$ and $B$ and is stable under local basis rotations, which can be regarded as free operations. Since $\mathcal{M}$ can be directly related to the flatness of the entanglement spectrum, a non-zero value of $\mathcal{M}$ directly implies that the state possesses both entanglement and magic [43]. For any faithful magic measure $M$, $\mathcal{M}$ would further be guaranteed to be non-negative. Moreover, if $M$ is additive, i.e., $M(\left|\psi\right\rangle\otimes\left|\phi\right\rangle)=M(\left|\psi\right\rangle)+M(\left|\phi\right\rangle)$, then the NM is subadditive. Specifically, for two bipartite pure states $\left|\Psi_{A_{1}B_{1}}\right\rangle$ and $\left|\Phi_{A_{2}B_{2}}\right\rangle$, we have $\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle)\geq\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)$, where $A=A_{1}\cup A_{2}$ and $B=B_{1}\cup B_{2}$. The proof can be found in [44]. In summary, for any additive and faithful magic measure, the NM is stable, subadditive, symmetric and vanishes for both product states and stabilizer states. Given these desirable properties and clear physical significance, $\mathcal{M}$ serves as an effective measure to characterize the novel resources of a quantum state that account for both magic and entanglement.
Figure 1: Definition of NM and two-qubit NM. (a) An illustration of how NM is defined. (b) A comparison of NM $\mathcal{M}$, von Neumann entropy $S$, and mutual magic $I$ for the state $\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$. Notice that $I$ is negative. As $\theta$ approaches 00 and $\pi/2$, the state becomes a product state, while at $\theta=\pi/4$, it is a stabilizer state. The total magic $M$ is also computed after applying a T-gate to the first qubit, where $\mathcal{M}$ remains unchanged. (c) The probability distribution of $\mathcal{M}$ for random pure states, with the red line representing the analytical result and the histogram depicting the numerical computation. (d) A comparison of $\mathcal{M}$ and logarithmic negativity $E_{N}$ for the Werner state. The total magic $M$ after applying a T-gate on the first qubit is also calculated. For the Werner state, the total magic $M$ always equals to the mutual magic $I$.
Although various magic measures satisfy faithfulness and additivity, we select $M$ to be the second stabilizer Rényi entropy (SRE) in the following, as it is well defined for qubits and avoids any minimization procedure in its calculation [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. It is given by:
$M(\rho)=-\text{log}\left(\frac{\sum_{P\in\mathbb{P}_{N}}|\text{Tr}(\rho P)|^{4}}{\sum_{P\in\mathbb{P}_{N}}|\text{Tr}(\rho P)|^{2}}\right),$ (2)
which for pure states reduces to $M(\left|\psi\right\rangle)=-\text{log}\sum_{P\in\mathbb{P}_{N}}(1/2^{N})|\left\langle\psi|P|\psi\right\rangle|^{4}$, where $N$ is the number of qubits and $\mathbb{P}_{N}$ represents the set of $N$-fold tensor products of Pauli matrices, denoted as $\sigma_{\mu}$ with $\mu=0,x,y,z$. It is also instructive to compare $\mathcal{M}$ with the mutual magic introduced in previous work, which is defined as $I_{AB}(\left|\Psi_{AB}\right\rangle)=M(\left|\Psi_{AB}\right\rangle)-M(\rho_{A})-M(\rho_{B})$, where $\rho_{A}$ and $\rho_{B}$ are reduced states of subsystems $A$ and $B$, respectively [55, 58, 57]. While the mutual magic captures certain aspects of the non-locality of a state’s magic, it lacks stability under local unitary transformations [61, 62], can take negative values in some cases, and generally does not align with the NM measure $\mathcal{M}$, as we will demonstrate later.
Two-qubit NM. In general, computing $\mathcal{M}$ is highly complicated because it involves minimizing over the full spectrum of local unitary transformations. However, for the two-qubit case, we demonstrate that an analytic expression for $\mathcal{M}$ can be derived. We first prove the following theorem:
Theorem: For a two-qubit pure state of the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$, $\mathcal{M}(\left|\psi\right\rangle)=M(\left|\psi\right\rangle)=\text{log}(\frac{8}{7+\text{cos}(8\theta)})$.
The proof is provided in [44], where the key idea is that the Pauli basis representation of this specific state is in a canonical form, with the reduced density matrices of both qubits exhibiting fully polarized Bloch vectors, and the full state having a diagonal correlation matrix [63, 64]. The following important corollary arises:
Corollary: For any two-qubit pure state $\left|\Psi\right\rangle$ with the bipartite entanglement spectrum $\{\text{cos}^{2}(\theta),\text{sin}^{2}(\theta)\}$, $\mathcal{M}(\left|\Psi\right\rangle)=\text{log}(\frac{8}{7+\text{cos}(8\theta)})$.
This holds because any two-qubit pure state can be transformed into the form $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$ by applying an appropriate local basis rotation, which can be determined via singular value decomposition. As a result, the NM is directly related to the entanglement spectrum and can be efficiently computed. It is worth noting that NM has previously been understood to be approximated by quantities derived from the entanglement spectrum [35]. Our results show that this approximation is, in fact, exact for the two-qubit case. A direct implication is that any two-qubit pure state can have at most $\text{log}(4/3)$ amount of NM, which is the maximal magic a single qubit can possess. Additionally, it is noteworthy that, when combined with the subadditivity of $\mathcal{M}$, if a many-body state can be decomposed into product states of multiple two-qubit states of the above form, an upper bound for its NM can be easily determined.
We compute $\mathcal{M}(\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle)$ for different values of $\theta$ and compare it with the von Neumann entropy $S$ and the mutual magic $I$. As shown in Fig. 1(b), $\mathcal{M}$ indeed vanishes for both stabilizer states and product states, while $S$ solely quantifies the entanglement of the state. Although the mutual magic $I$ exhibits a similar trend to $\mathcal{M}$, it take negative values, and its peak differs from that of $\mathcal{M}$. Notably, while for this specific form of state, $\mathcal{M}$ is equal to the total magic $M$, differences arise when a local gate is applied to either qubit. For instance, when a T-gate is applied to the first qubit, the NM remains unchanged, but the total magic $M$ changes significantly, displaying a non-zero value at $\theta=\pi/4$.
Subsequently, we investigate the typical amount of NM in a random two-qubit pure state. We generate random pure states via a Haar-random two-qubit unitary gate and examine the distribution of $\mathcal{M}$ across these states. The results, shown in Fig. 1(c), indicate that the state most commonly hosts the maximal $\mathcal{M}$. While the analytical expression for this probability distribution can be derived [44], a closed-form expression for the average value of $\mathcal{M}$, analogous to that for entanglement [65, 66], remains unclear. Nevertheless, numerical results suggest that the average $\mathcal{M}$ is approximately $0.192$.
Mixed state generalization. For mixed states, the NM can be defined analogously to Eq. (1), with $\left|\Psi_{AB}\right\rangle$ replaced by the density matrix $\rho_{AB}$. Under this definition, the NM retains the properties of stability, symmetry and subadditivity for any additive magic measure. By continuing to employ the second SRE as the magic measure, which is faithful when mixed stabilizer state are regarded as resource-free, the NM also vanishes for mixed stabilizer states [45, 67]. A crucial difference, however, is that the NM does not vanish for certain separable states. This difference arises because, even for a single qubit mixed state, such as $\rho=\frac{1}{2}(\sigma_{0}+\frac{1}{2}\sigma_{x})$, no unitary transformation can bring it back to a mixed stabilizer state. We also explore other possible magic measures for mixed state, while the conclusion remains the same [44].
Focusing once again on the two-qubit case, it follows directly from the proof of the previous theorem that if a two-qubit mixed state $\rho$ has a Pauli basis representation in canonical form, we would have $\mathcal{M}(\rho)=M(\rho)$. However, it is important to note that not all mixed state can be transformed into the canonical form using only local unitary transformations [63]. In such cases, numerical optimization is inevitable, as examples will be shown in the next section.
Consider the Werner state as an analytically tractable example, which is defined as $\rho_{W}=x\left|\psi_{-}\right\rangle\left\langle\psi_{-}\right|+(1-x)\mathbb{I}/4$, with $|\psi_{-}\rangle\equiv(|01\rangle-|10\rangle)/\sqrt{2}$ [68]. This state has vanishing Bloch vectors for each qubit and a diagonal correlation matrix $T$ with diagonal entries given by $-x/4$. We calculate NM and logarithmic negativity $E_{N}$, a measure of separability, with the results shown in Fig. 1(d). Although the Werner state remains separable for $x<1/3$, the NM takes on a finite value for small $x$. As $x$ approaches $1$, the state becomes more entangled, yet the NM approaches zero as the state becomes a stabilizer state. The difference between $\mathcal{M}$ and total magic $M$ again becomes evident if we further apply a T-gate on the first qubit. While the NM remain unaffected, the total magic $M$ would change significantly.
Figure 2: Results for the TFIM with system size $L=128$. The ground states are obtained using density matrix renormalization group method, with a maximum bond dimension $\chi=500$. (a) Two-point NM for different $h$. The inset displays a power-law fit with a constant shift for $h=1.0$. (b-d) MINM for measurements along the $x-$, $y-$ and $z-$axes, respectively. The insets show the power-law fit for $h=1.0$.
NM in many-body state. Building on our improved understanding of the NM in the two-qubit case, we now shift our focus to the scaling behavior of the two-point NM in many-body quantum states, defined as $\mathcal{M}(\rho_{1,r})$, where $\rho_{1,r}$ represents the reduced density matrix for the first qubit and the $r$-th qubit. For systems with translational symmetry and periodic boundary conditions, the two-point NM depends solely on the relative separation $r$. The scaling behavior provides valuable information on the properties of many-body states, especially for critical states. We also explore the measurement-induced non-local magic (MINM), defined as ~$\widetilde{\mathcal{M}}(r)\equiv\sum_{i}p_{i}\mathcal{M}(\left|\Psi_{1,r}^{i}\right\rangle)$, where $p_{i}$ is the probability of obtaining the outcome $i$ from projective measurements performed on all qubits except the first and the $r$-th qubits, and $\left|\Psi_{1,r}^{i}\right\rangle$ represents the corresponding post-measurement state. Numerical details are provided in Supplemental Material [44].
We first consider the TFIM, with Hamiltonian given by $H_{\text{TFIM}}=-\sum_{i}\sigma_{x}^{i}\sigma_{x}^{i+1}-h\sum_{i}\sigma_{z}^{i}$. The ground state is paramagnetic for $h>1$ and ferromagnetic for $h<1$, with a quantum phase transition occurring at $h=1$. Due to the $\mathbb{Z}_{2}$ symmetry $\prod_{i}\sigma_{x}^{i}$, the Pauli basis representation of $\rho_{1,r}$ is constrained to be in the canonical form, enabling direct computation of the NM. As shown in Fig. 2(a), the NM exhibits a power-law decay with an exponent of around $0.5$, close to that of mutual information. In Ref. [44], we show that their decay exponents are analytically identical for sufficiently large $r$. Away from criticality, NM decays exponentially to a constant value, distinguishing it from critical states.
Next, we examine the MINM for different measurement directions. As illustrated in Fig. 2(b-d), the MINM exhibits a power-law decay for critical states, with different decay exponents depending on the measurement axis. The MINM under $y-$ and $z-$axes measurements decays more slowly than the pre-measurement behavior, but still faster than the slowest decaying correlation prior to measurement, namely the $\left\langle\sigma_{x}^{1}\sigma_{x}^{r}\right\rangle$, which decays with exponent $1/4$ [69]. For non-critical states, the MINM either becomes constant or decays exponentially.
Figure 3: Results for critial MHC with $L=48$, with a maximum bond dimension $\chi=1000$. (a) Haar-random unitaries are applied in a brick-wall manner, with a measurement rate of $p=0.17$ [40]. For the MINM, all qubits except the first and the $r$-th qubits are projectively measured along the $z$-axis. (b) A comparison between averaged two-point NM $\overline{\mathcal{M}}$ and total magic $\overline{M}$. Each data point is averaged over 3000 trajectories. The inset shows the power-law fit with a constant shift for $\overline{\mathcal{M}}$. (c) The MINM and the average total magic after measurement, with each data point averaged over $4\times 10^{4}$ trajectories. (d) A comparison between the scaling behavior of the MINM and the mutual information $I_{E}$ prior measurement. The decay exponents for $\widetilde{\mathcal{M}}$ and $I_{E}$ are around $0.76$ and $3.31$, respectively.
We then consider the MHC at a critical measurement rate of $p=0.17$. The circuit structure is shown in Fig. 3(a). We consider the average NM defined as ¯$\overline{\mathcal{M}}=\mathbb{E}_{U}\mathbb{E}_{m}[\mathcal{M}]$, where $\mathbb{E}_{U}$ and $\mathbb{E}_{m}$ represent the averaging over unitary gates and measurement outcomes, respectively. Since the model is structureless, with only locality as a constraint, $\rho_{1,r}$ does not necessarily have a canonical Pauli basis representation, and the NM must be computed by numerical optimization. As shown in Fig. 3(b), NM is significantly lower than the total magic and also feature a power-law decay. We then consider the MINM and compare it with the average total magic for post-measurement states, as presented in Fig. 3(c). The measurement direction is inessential, and we assume it to be along the $z$-axis. It is evident that a greater amount of magic can be removed by local unitary transformations as the distance increases. Furthermore, we observe that the decay exponent of MINM is smaller than half of that of mutual information prior to measurement, as shown in Fig. 3(d). Since the slowest correlation a state can maintain must decay at an exponent at least half that of mutual information [70], this suggests that the measurement induces NM with a longer range than any pre-existing correlations. This phenomenon is akin to entanglement swapping, where measurements induce long-range entanglement between previously uncorrelated qubits. We refer to this effect as “magic swapping”. Since it has been established that $\mathcal{M}$ is upper-bounded by entanglement [35], and that measurement-induced entanglement cannot decay more slowly than correlations in sign-free states [71], it follows that magic swapping can only occur in states possessing sign structures.
Discussions. As we focus on the NM for qubit systems, a natural extension is to investigate systems with higher-dimensional local Hilbert spaces, such as qutrits. In systems where the local Hilbert space dimension is an odd prime, mana serves as an alternative, appropriate magic measure for both pure and mixed states [72, 73, 74, 75]. Notably, like SRE, mana does not necessitate any minimization procedure. However, it remains an open question whether an analytical expression can be derived for the two-qutrit case. It would also be interesting to investigate the consistency between the results obtained using SRE and mana. We leave this to future work.
While NM can be expressed analytically for two-qubit pure states, making it computationally straightforward and feasible for experimental measurements, the situation becomes more complex for general multi-qubit states. In this case, computing the NM presents an optimization problem involving $O(N^{2})$ parameters, which becomes increasingly difficult as the number of qubits grows. One approach to address this challenge is to employ automatic differentiation techniques, similar to those used in simulating the variational quantum eigensolver algorithm, with the target function being the total magic rather than the energy [76]. Scaling up the calculation of NM would facilitate the classification of states based on whether they exhibit volume-law or area-law NM, analogous to the classification of states in terms of entanglement. We conjecture that only states that exhibit both volume-law entanglement entropy and volume-law magic can sustain a volume-law NM.
Additionally, it is noteworthy that a related measure, non-stabilizerness entanglement entropy (NsEE), has recently been introduced [77, 78]. In this formulation, free operations are defined as global Clifford gates instead of local unitary gates, and the quantity calculated is the entanglement entropy of the total state, rather than its magic. While NsEE and NM aim to characterize the “quantumness” of a system from distinct perspectives, further investigation into the relationship between these two measures would be a valuable direction for future research.
Acknowledgements.
Acknowledgments. We thank Xiaoyu Dong for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01.
Supplemental Material for “Quantum Non-Local Magic”
Dongheng Qian1, 2 and Jing Wang1, 2, 3, ${}^{1,\,2,\,3,\,4,\,*}$
1State Key Laboratory of Surface Physics and Department of Physics, Fudan University, Shanghai 200433, China
2Shanghai Research Center for Quantum Sciences, Shanghai 201315, China
3Institute for Nanoelectronic Devices and Quantum Computing, Fudan University, Shanghai 200433, China
4Hefei National Laboratory, Hefei 230088, China
S1 Analytic results
S1.1 Proof of subadditivity
In this section, we prove that the non-local magic (NM), defined as $\mathcal{M}_{AB}(\left|\Psi_{AB}\right\rangle)=\min_{U_{A}\otimes U_{B}}M(U_{A}\otimes U_{B}\left|\Psi_{AB}\right\rangle)$, is subadditive if the magic measure $M$ is additive. By definition, we have:
$\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)=\min_{{U_{A}\otimes U_{B}}}M(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)=\min_{U_{A}\otimes U_{B}}[M(\left|\Psi_{A_{1}B_{1}}\right\rangle)+M(\left|\Phi_{A_{2}B_{2}}\right\rangle)].$ (S1)
The second equation follows directly from the additivity of $M$ and we denote $A=A_{1}+A_{2}$ and $B=B_{1}+B_{2}$. Since unitaries of the form $U_{A_{1}}\otimes U_{A_{2}}\otimes U_{B_{1}}\otimes U_{B_{2}}$ form a subset of the set of unitaries of the form $U_{A}\otimes U_{B}$, we further have:
$\mathcal{M}_{AB}(\left|\Psi_{A_{1}B_{1}}\right\rangle\otimes\left|\Phi_{A_{2}B_{2}}\right\rangle)\leq\min_{U_{A_{1}}\otimes U_{B_{1}}}M(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\min_{U_{A_{2}}\otimes U_{B_{2}}}M(\left|\Phi_{A_{2}B_{2}}\right\rangle)=\mathcal{M}_{A_{1}B_{1}}(\left|\Psi_{A_{1}B_{1}}\right\rangle)+\mathcal{M}_{A_{2}B_{2}}(\left|\Phi_{A_{2}B_{2}}\right\rangle).$ (S2)
This establishes the subadditivity of $\mathcal{M}$.
Building on this subadditivity, it is straightforward to establish an upper bound for the NM of a many-body state, provided that it can be brought into product states of two-qubit pure states in canonical form via local unitary transformations. It has been shown that such transformations can be approximately realized using the Petz recovery map for Haar-random states [1]. However, the exact determination of such a transformation remains an open problem and is a subject worth further exploration.
S1.2 Proof of theorem
In this section, we provide a detailed proof of the theorem introduced in the main text. Any two-qubit state $\rho$ consisting of qubits $A$ and $B$ can be expressed in Pauli basis as:
$\rho=\frac{1}{4}\sum_{\mu,\nu}R_{\mu\nu}\sigma_{\mu}\sigma_{\nu},$ (S3)
where $\mu,\nu\in\{0,x,y,z\}$, and the coefficients $R$ are determined by $R_{\mu\nu}=\text{Tr}(\rho\sigma_{\mu}\otimes\sigma_{\nu})$ [2, 3]. The matrix $R$ takes a block structure of the form:
$R=\begin{pmatrix}1&\textbf{a}^{T}\\ \textbf{b}&T\end{pmatrix}$ (S4)
where a and b are the Bloch vectors corresponding to the reduced density matrices of qubit $A$ and $B$, respectively, and $T$ represents the correlation matrix. In this representation, the second stabilizer Rényi entropy (SRE) of the state $\rho$ is concisely expressed as:
$M(\rho)=-\text{log}\left(\sum_{\mu\nu}{|R_{\mu\nu}|^{4}}\right)-S_{2}(\rho),$ (S5)
where $S_{2}(\rho)$ denotes the 2-Rényi entropy of the state $\rho$, which is invariant under any unitary transformation applied to $\rho$. Therefore, the calculation of NM $\mathcal{M}$ reduces to finding the appropriate local unitary transformation that maximizes the sum $\sum_{\mu\nu}{|R_{\mu\nu}|^{4}}$.
Any local unitary transformation of the form $U=U_{A}\otimes U_{B}$ corresponds to a basis rotation, resulting in the transformations $\rightarrow O_{A}\textbf{a}$ a, $\rightarrow O_{B}\textbf{b}$ b and $T\rightarrow O_{A}TO_{B}^{T}$, where $O_{A}$ and $O_{B}$ are orthogonal matrices corresponding to rotations in qubits $A$ and $B$, respectively. Since $O_{A}$ is orthogonal, it preserves the norm of the Bloch vector a, and we have
$\sum_{i}\left|a_{i}\right|^{4}\leq\sum_{i}\left|a_{i}\right|^{2}=\text{const},$ (S6)
with equality holding if and only if a is either fully polarized or the zero vector. The same statement applies to b. Furthermore, if the matrix $T$ is diagonal, i.e., $T=\text{diag}\{d_{1},d_{2},d_{3}\}$, the effect of $O_{A}$ on the $j$-th column of $T$ is governed by the following inequality:
$\sum_{i=1}^{3}\left|T^{\prime}_{i,j}\right|^{4}=\sum_{i=1}^{3}\left|d_{j}O_{A,ij}\right|^{4}=\left|d_{j}\right|^{4}\sum_{i=1}^{3}\left|O_{A,ij}\right|^{4}\leq\left|d_{j}\right|^{4}\sum_{i=1}^{3}\left|O_{A,ij}\right|^{2}=\left|d_{j}\right|^{4},$ (S7)
where the final equality follows from the orthogonality of $O_{A}$. A similar argument holds for the action of $O_{B}$. This indicates that a diagonal matrix $T$ maximizes the fourth power sum of its elements under any local unitary transformation. Combining these results, we conclude that any local unitary transformation applied to a two-qubit state with fully polarized Bloch vectors for both qubits and a diagonal correlation matrix increases its magic. We call such a state to be in the canonical form of its Pauli basis representation. It follows directly that if a state $\rho$ is in canonical form, then $\mathcal{M}(\rho)=M(\rho)$.
For the specific state $\left|\psi\right\rangle=\text{cos}(\theta)\left|00\right\rangle+\text{sin}(\theta)\left|11\right\rangle$, the corresponding $R$ matrix is:
$R=\begin{pmatrix}1&0&0&\text{cos}(2\theta)\\ 0&\text{sin}(2\theta)&0&0\\ 0&0&-\text{sin}(2\theta)&0\\ \text{cos}(2\theta)&0&0&1\end{pmatrix},$ (S8)
which clearly exhibits a canonical form. This completes the proof of the theorem stated in the main text.
It is worth noting that while a local unitary transformation always exists that brings a two-qubit pure state into canonical form, this is not generally true for a mixed state. However, a unique normal form exists for any two-qubit mixed state, where the correlation matrix $T$ is diagonal, although the Bloch vectors of the qubits need not be fully polarized [4].
S1.3 Probability distribution
In this section, we derive the analytic expression for the probability distribution of NM for the two-qubit Haar random state. It is well known that the Haar random state has an entanglement spectrum $\{\text{cos}(\theta),\text{sin}(\theta)\}$, with the associated probability distribution given by $P(x)=3(1-2x)^{2}$, where we define $x\equiv\text{cos}^{2}(\theta)$ [5, 6]. From the corollary stated in the main text, the NM for a state with such an entanglement spectrum is expressed as:
$\mathcal{M}(x)=\text{log}\left(\frac{8}{7+\text{cos}(8\theta)}\right)=\text{log}\left(\frac{1}{1+4x(x-1)(1-2x)^{2}}\right).$ (S9)
Our goal is to determine the probability distribution of $y\equiv\mathcal{M}(x)$, where $x$ is a random variable following the distribution $P(x)$. The probability distribution of $y$ is then given by:
$P(y)=\sum_{x:\mathcal{M}(x)=y}\frac{P(x)}{\mathcal{M}^{\prime}(x)},$ (S10)
where $\mathcal{M}’(x)$ is the derivative of $\mathcal{M}(x)$ with respect to $x$, which can be explicitly calculated as:
$\mathcal{M}^{\prime}(x)=-\frac{4(2x-1)(1+8x(x-1))}{1+4(2x-1)^{2}(x-1)x}.$ (S11)
Furthermore, the solution to the equation $y=\mathcal{M}(x)$ is:
$x=\frac{1}{2}\pm\frac{1}{4}\sqrt{2\pm 2e^{-y/2}z}.$ (S12)
For simplicity, we introduce the notation $z\equiv\sqrt{4-3e^{y}}$. Substituting Eq. (S11) and Eq. (S12) into Eq. (S10), we finally obtain the probability distribution of $y$ as:
$P(y)=\frac{3e^{-y/2}((e^{y/2}+z)\sqrt{1-e^{-y/2}z}+(e^{y/2}-z)\sqrt{1+e^{-y/2}z})}{4z\sqrt{2(e^{y}-1)}}.$ (S13)
[…]
S1.4 Decay exponent for TFIM
In this section, we demonstrate that the decay exponent of NM in the critical ground state of TFIM is identical to that of mutual information. We begin by deriving a general result for the decay exponent of the two-point total magic. At the critical point, all connected correlation functions of the form $\mathcal{C}_{\mu\nu}=\text{Tr}(\rho_{1,r}\sigma_{\mu}^{1}\sigma_{\nu}^{r})-\text{Tr}(\rho_{1}\sigma_{\mu}^{1})\text{Tr}(\rho_{r}\sigma_{\nu}^{r})$ obey a power-law decay, i.e., $\mathcal{C}_{\mu\nu}\sim r^{-\alpha_{\mu\nu}}$, where $\mu,\nu$ takes values in $\{0,x,y,z\}$. For the correlations that vanish, we assume that the corresponding decay exponent $\alpha$ is infinite. Thus, we can write the correlation function as:
$\text{Tr}(\rho_{1,r}\sigma_{\mu}^{1}\sigma_{\nu}^{r})\sim r^{-\alpha_{\mu\nu}}+c_{\mu\nu},$ (S14)
where $c_{\mu\nu}\equiv\text{Tr}(\rho_{1}\sigma_{\mu}^{1})\text{Tr}(\rho_{r}\sigma_{\nu}^{r})$ is independent of $r$ due to translational invariance. Using this, the total magic of the state can be expressed as:
$M(\rho_{1,r})\sim\text{log}\left(\frac{\sum_{\mu\nu}(r^{-\alpha_{\mu\nu}}+c_{\mu\nu})^{2}}{\sum_{\mu\nu}(r^{-\alpha_{\mu\nu}}+c_{\mu\nu})^{4}}\right)\sim\text{log}\left(\frac{c_{1}+\sum_{\mu\nu}2c_{\mu\nu}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}r^{-2\alpha_{\mu\nu}}}{c_{2}+\sum_{\mu\nu}4c_{\mu\nu}^{3}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}6c_{\mu\nu}^{2}r^{-2\alpha_{\mu\nu}}+\dots}\right),$ (S15)
where $c_{1}$ and $c_{2}$ represent the summations of the constant terms that are independent of $r$ in the numerator and denominator, respectively. The ellipses indicate terms that decay with higher exponents. In the large $r$ limit, we can further approximate as:
$M(\rho_{1,r})\sim\text{log}\left(\frac{c_{1}}{c_{2}}\right)+\left[1+\frac{1}{c_{1}}(\sum_{\mu\nu}2c_{\mu\nu}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}r^{-2\alpha_{\mu\nu}})\right]\left[1-\frac{1}{c_{2}}(\sum_{\mu\nu}4c_{\mu\nu}^{3}r^{-\alpha_{\mu\nu}}+\sum_{\mu\nu}6c_{\mu\nu}^{2}r^{-2\alpha_{\mu\nu}}+\dots)\right],$ (S16)
where $(1+x)^{-1}\sim(1-x)$ and $\text{log}(1+x)\sim x$ for small $x$ are used. Denoting $\alpha^{*}\equiv\min_{\mu\nu}\alpha_{\mu\nu}$ and $\overline{\alpha}\equiv\min_{\mu\nu}\{\alpha_{\mu\nu};c_{\mu\nu}\neq 0\}$, it follows that the in the large $r$ limit, the total magic $M$ is governed by:
$M(\rho_{1,r})\sim\text{const}+r^{-2\alpha^{*}}+r^{-\overline{\alpha}}.$ (S17)
Thus, the decay exponent of the total magic is given by:
$\alpha_{M}=\text{min}(2\alpha^{*},\overline{\alpha}).$ (S18)
This conclusion holds generally for any particular critical state.
At the critical point of the TFIM, we have $\alpha^{*}=\alpha_{xx}=1/4$ and $\overline{\alpha}=\alpha_{zz}=2$ [7], which results in $\alpha_{M}=1/2$, identical to the decay exponent of the mutual information. Furthermore, since the ground state is in canonical form, the NM is equivalent to the total magic of the two-qubit state. This completes the proof.
S2 Numerical results
S2.1 Numerical details
In this section, we provide numerical details for the simulations conducted in the main text. The ground state of the TFIM is obtained using the density matrix renormalization group algorithm using the ITensor package [8, 9], with the maximum bond dimension $\chi=500$ and truncation cutoff $\epsilon=10^{-12}$. For the monitored Haar-random unitary circuit at the critical point, the time evolution was performed by the time-evolving block decimation algorithm, with the maximum bond dimension $\chi=1000$ and truncation cutoff $\epsilon=10^{-6}$.
When calculating NM through numerical optimization over all local unitary transformations, we parameterize the local unitary on a single qubit as $U=\begin{pmatrix}\text{cos}(\theta/2)&-e^{i\lambda}\text{sin}(\theta/2)\\ e^{i\phi}\text{sin}(\theta/2)&e^{i(\phi+\lambda)}\text{cos}(\theta/2)\end{pmatrix}$, where $\phi\in[0,2\pi]$, $\theta\in[0,\pi]$ and $\lambda\in[0,2\pi]$. The optimization process employed the Nelder-Mead algorithm, as implemented in the Optim package [10]. To mitigate the risk of converging to a local minimum, we initiated the optimization from 100 distinct starting points and selected the minimal value achieved.
S2.2 Separable mixed states host non-zero NM
In this section, we demonstrate that for any choice of magic measure for mixed states, certain separable states will always exhibit a non-zero NM. A separable state is defined as the convex hull of pure product states [11]:
$\rho=\sum_{i}p_{i}\rho_{iA}\otimes\rho_{iB}.$ (S19)
Within the framework of resource theory, there are two inequivalent ways to define mixed stabilizer states, which we denote as $\text{STAB}_{0}$ and STAB, following the notation in Ref. [12]. $\text{STAB}_{0}$ is defined as the set of states $\rho=\frac{1}{2^{N}}\sum_{P\in G}P$, where $G$ is a group of commuting Pauli operators $P$ [13], while STAB is defined as the convex hull of pure stabilizer states:
$\text{STAB}:=\{\rho|\rho=\sum_{i}p_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|\},$ (S20)
where $\psi_{i}$ are pure stabilizer states. When STAB is regarded as the set of free states, it specifically captures distillable non-Clifford resources. On the other hand, regarding $\text{STAB}_{0}$ as free states, it also incorporates the non-flat probability distributions of mixed states as resources. In the second SRE used in the main text, $\text{STAB}_{0}$ is treated as the set of free states, while other magic measures, such as the robustness of magic (RoM) [14, 15, 16] and mana [17], consider STAB as the set of free states. It is also worth mentioning that the SRE has been extended by a convex-roof construction to provide an alternative measure in the latter case [18].
As outlined in the main text, unitary transformations cannot transform even a single-qubit state into one with a flat probability distribution, i.e., $\min_{U}M(U\rho U^{\dagger})$ is not necessarily zero. This directly implies that for a separable state $\rho_{A}\otimes\rho_{B}$, we have:
$\mathcal{M}(\rho_{A}\otimes\rho_{B})=\min_{U_{A}}M(\rho_{A})+\min_{U_{B}}M(\rho_{B})\neq 0.$ (S21)
This scenario changes when STAB is considered as the set of free states, as any single-qubit density matrix can always be diagonalized, thus ensuring that it lies within STAB. However, we are unaware of any proof guaranteeing that every separable multi-qubit state can be transformed into a state belonging to STAB using only local unitary transformations, and we conjecture that this is not the case. To provide numerical evidence, consider the separable state:
$\rho_{0}=\frac{1}{2}\left|\phi_{0}\phi_{0}\right\rangle\left\langle\phi_{0}\phi_{0}\right|+\frac{1}{2}\left|00\right\rangle\left\langle 00\right|.$ (S22)
Here, $\left|\phi_{0}\right\rangle=\text{cos}(\pi/8)\left|0\right\rangle+\text{sin}(\pi/8)\left|1\right\rangle$. Using RoM as the magic measure, which is defined as $R(\rho)=\min_{x}\{\sum_{i}|x_{i}|;\rho=\sum_{i}x_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|\}-1$, we numerically minimize the magic over all possible local unitary transformations and find that $\mathcal{M}_{R}(\rho_{0})\approx 0.0703$, which is non-zero. This indicates that even when using magic measures that treat STAB as the set of free states, certain separable states can still exhibit a non-zero NM.